I don't understand this math and the book doesn't give a good explanation of how to solve the problem.
1/6t = 18

x+8/4= 6

Answer by checkley77(12569)

1/6t = 18
t/6=18
t=6*18
t=108 ans.
-------------------------------
x+8/4=6
x=6-8/4
x=6-2
x=4 ans.